# MemeCLIP: Leveraging CLIP Representations for Multimodal Meme Classification

### v1

@@ Abstract
The complexity of text-embedded images presents a formidable challenge in machine learning given the need for multimodal understanding of the multiple aspects of expression conveyed in them. While previous research in multimodal analysis has primarily focused on singular aspects such as hate speech and its subclasses, our study expands the focus to encompass multiple aspects of linguistics: hate, target, stance, and humor detection. We introduce a novel dataset PrideMM comprising text-embedded images associated with the LGBTQ+ Pride movement, thereby addressing a serious gap in existing resources. We conduct extensive experimentation on PrideMM by using unimodal and multimodal baseline methods to establish benchmarks for each task. Additionally, we propose a novel framework MemeCLIP for efficient downstream learning while preserving the knowledge of the pre-trained CLIP model. The results of our experiments show that MemeCLIP achieves superior performance compared to previously proposed frameworks on two real-world datasets. We further compare the performance of MemeCLIP and zero-shot GPT-4 on the hate classification task. Finally, we discuss the shortcomings of our model by qualitatively analyzing misclassified samples. Our code and dataset are publicly available at: https://github.com/SiddhantBikram/MemeCLIP.
@@
Figure 1: Examples of text-embedded images labeled for all aspect labels in the PrideMM dataset. The labels are in the form of {Hate, Target, Stance, Humor}. For Hate, {0, 1} correspond to No Hate and Hate respectively. For Target, {0, 1, 2, 3} correspond to hate targeted towards Undirected, Individual, Community, and Organization respectively. For Stance, {0, 1, 2} correspond to Neutral, Support, and Oppose respectively. For Humor, {0, 1} correspond to No Humor and Humor respectively.
@@ 1 Introduction
However, with opinions being expressed freely, hate speech becomes prevalent, often targeted towards individuals, organizations, and even marginalized communities Thapa et al. (2023), targeting them with vitriol and prejudice Lingiardi et al. (2020); Imperato et al. (2023). Particularly, the LGBTQ+ movement stands as a prominent exemplar of online discourse, where memes serve as vehicles of both solidarity and resistance, reflecting the multifaceted dynamics of attitudes and perceptions within the community and beyond Gal et al. (2016). In this context, the distinction between humor and harm becomes blurred, as memes straddle the line between satire and offense, challenging researchers and platforms alike to navigate the complexities of online content moderation Langvardt (2017). Previous attempts that endeavored to suppress such content have resulted in the discriminative suppression of all LGBTQ+ content Griffin (2022, 2024), which can harm the awareness and acceptance of this community. Thus, understanding the nuances of hate speech, opinions, and intended humor within memes is paramount for fostering an inclusive digital environment and combating online discrimination.
@@
The analysis of text-embedded images is particularly challenging given the need for contextual understanding and the prevalence of ambiguity and subjectivity in them Sherratt (2022). The multi-aspect nature of our dataset provides a more holistic view of the diverse themes usually expressed through memes. Samples of images and their annotations are illustrated in Figure 1. Through PrideMM, we aim to cultivate a more profound understanding of interaction dynamics through memes on social media and facilitate the development of multimodal content moderation methods in order to make the internet a safer space. We implement a range of baseline and state-of-the-art hate speech detection models to establish benchmarks for each task of PrideMM.
We further propose MemeCLIP, a novel framework that leverages the knowledge of CLIP (Contrastive Language-Image Pre-Training) Radford et al. (2021) for multi-aspect meme classification. We employ linear layers to effectively disentangle image and text representations in CLIP’s multimodal embedding space. We utilize lightweight Feature Adapters to preserve the prior knowledge of CLIP and adapt its embedding spaces to the meme classification task while avoiding overfitting on smaller datasets. We further integrate a cosine classifier alongside Semantic-Aware initialization Shi et al. (2023) to make MemeCLIP more robust to the imbalanced class labels that may exist in datasets such as PrideMM and HarMeme Pramanick et al. (2021a) that are representative of real-world data distributions. Distinct from previous multimodal meme classification frameworks, MemeCLIP is trained end-to-end in a single step and does not rely on extraneous models to create augmented data. Our main contributions can be summarized as follows:
@@
  • We benchmark PrideMM by using various unimodal and multimodal methods including previously proposed multimodal frameworks for hate speech detection.
@@ 2.1 Multimodal Datasets
Multimodal image-text analysis has seen significant strides in recent years, owing to its widespread popularity and high availability across social media platforms. Qu et al. (2022) introduced the DisinfoMeme dataset that contains memes from three different contexts: the COVID-19 pandemic, the Black Lives Matter (BLM) movement, and Veganism, labeled for disinformation identification. In an effort to discern humor often expressed in memes, Tanaka et al. (2022) created a humor detection dataset by proposing a refined process for extracting memes devoid of interpersonal influence on the perception of humor. With the increasing need for hate speech and offensive content detection in social media, multimodal data analysis for hate speech has seen a particular surge. The HMC dataset Kiela et al. (2020) was released as part of the Hateful Meme Challenge and contains synthetic memes designed to convey contrastive meanings from image and text modalities. Hateful memes in the dataset target religion, race, disability, and sex. Suryawanshi et al. (2020) employed extensive annotation guidelines to create the MultiOFF dataset, consisting of memes collected from Reddit, Facebook(FB), Twitter, and Instagram for offensive content detection. The Harm-C Pramanick et al. (2021a) and Harm-P Pramanick et al. (2021b) datasets comprise memes related to the COVID-19 pandemic and US politics respectively. Both datasets were annotated for three degrees of harmfulness and four subclasses for the targets of hate speech. Similarly, Bhandari et al. (2023) collected text-embedded images from social media platforms - Twitter, Facebook, and Reddit related to the Russia-Ukraine Conflict for hate speech detection, annotating samples for hate speech detection and target detection.
@@ 2.2 Multi-aspect Datasets
Online discourse on socio-political events is often imbued with a series of human emotions, which has led researchers to study the multiple aspects of linguistics expressed in them. Dacon et al. (2022) used comments collected from RedditBias Barikeri et al. (2021) related to LGBTQ+ individuals and annotated each comment for the presence of Toxicity, Severe Toxicity, Obscene, Threat, Insults, and Identity Attacks. Similarly, Gautam et al. (2020) curated a dataset of tweets related to the #MeToo movement in social media by annotating the tweets in five different aspects. Taking the studies of multiple aspects one step further, Ousidhoum et al. (2019) compiled an extensive multi-aspect Twitter dataset in English, French, and Arabic with each language dataset annotated for different aspects including hate and offensiveness. Table 1 provides a detailed comparison of the datasets used in the literature.
Multi-aspect data helps better understand the spectrum of human emotions that may be associated with social media interactions. Most multimodal datasets, while only focusing on a single aspect and its sub-classes, fail to encompass the complex dynamics of emotions expressed by the masses. Our work aims to address this gap by presenting a multimodal and multi-aspect dataset comprising three different aspects- hate detection, stance detection, and humor detection, and one subclass within hate- target detection, to enable more nuanced studies of multimodal meme data through computational methods.
@@ 2.3 Multimodal Frameworks
Recent developments in large vision-language models have incited a wave of research in methods to tackle hate speech in text-embedded images. MOMENTA Pramanick et al. (2021b) was one of the first frameworks proposed to incorporate CLIP’s vision and language encoders for multimodal hate speech classification. They further extract regions of interest from image data and named entities from text data and combine them with CLIP representations by using cross-modal attention fusion. Similarly, HateCLIPper Kumar and Nandakumar (2022) was proposed to better model cross-modal interactions between CLIP representations. Textual inversion Gal et al. (2022) has been used to integrate visual cues in the text representation space in frameworks such as ISSUES Burbi et al. (2023). Recent works make use of image caption models to extract text captions from images and learn a single language processing model Cao et al. (2023b, a). However, rather than relying on augmented data from extraneous models, our proposed framework MemeCLIP leverages the knowledge learned by CLIP’s encoders during its pre-training step to process the rich multimodal information inside each image. Additionally, we use Feature Adapters alongside residual connections to prevent overfitting as annotated datasets for multimodal meme classification generally lack a high number of samples. We further utilize a cosine classifier to make MemeCLIP more robust to imbalanced data classes, which is prevalent in multi-label tasks in this domain.
@@ 3 Dataset
In this section, we describe various aspects of our dataset including data collection, annotation guidelines, and dataset statistics. Our dataset comprises 5,063 text-embedded images that encompass memes, posters, and infographics relevant to the LGBTQ+ movement. We include images from 2020-2024 as this period saw an upsurge of social media content in this domain Oz et al. (2023). This also allows our dataset to represent contemporary social media interactions through memes. Note that by the term LGBTQ+, we refer to all gender identities and sexual orientations.
@@ 3.1 Data Collection
To maintain diversity in the dataset, we collected data from three popular social media platforms- Facebook, Twitter, and Reddit, through manual search and extraction. For Twitter, we used hashtags such as #lgbt, #pride, #trans, #transrights, #nonbinary, and #genderidentity to filter images related to LGBTQ+ discussions. For Facebook, we targeted groups that frequently discussed LGBTQ+ content. Similarly, for Reddit, we identified subreddits where discussion related to LGBTQ+ was more prominent. Further, to ensure the relevance and quality of the dataset, the data collection process was subject to filtering criteria. Detailed filtering criteria for our dataset can be found in Appendix A.2. Since different annotators may collect duplicates of the same image, we sequentially employed two image deduplication tools, dupeGuru [^1] https://github.com/arsenetar/dupeguru and difPy [^2] https://github.com/elisemercury/Duplicate-Image-Finder, to filter out duplicates and retain the highest quality images out of the batch of duplicates. We used the OCR application provided by Google Cloud Vision API [^3] https://cloud.google.com/vision/docs to extract textual data from the images. We removed non-alphanumeric elements such as special characters, hyperlinks, symbols, and non-English characters to reduce noisy text data and ensure data quality. Note that the text may occasionally contain unintentional noisy artifacts.
@@ 3.2 Data Annotation
We engaged five experienced annotators, well-versed in NLP and computational linguistics, to annotate data samples for PrideMM. The annotators had a prior understanding of the LGBTQ+ movement and meme dynamics on social media. We presented them with comprehensive annotation guidelines to ensure uniform and unbiased annotations and asked them to annotate each image separately for all four tasks. A 3-phase annotation schema was used to ensure accurate and consistent annotations. First, a dry run was conducted where the annotators were given 50 images each for annotation to evaluate the understanding of the annotation guidelines among the annotators. Second, a revision phase was conducted where the annotators were given another 200 images each and received a revised set of instructions based on the results of the first phase. Finally, in the consolidation phase, the annotators annotated 50 images while discussing and revising the annotation guidelines until a consensus was reached. These steps were taken to avoid misannotations and noisy labels in PrideMM as much as possible. We followed meticulously devised annotation guidelines to ensure consistency in the annotations. Every image in our dataset was annotated for three different aspects and one sub-class. Apart from the relationship between ’Hate’ and ’Hate Targets’, all aspects were annotated independently.
@@ 3.3 Annotation Guidelines
Hate Speech: This task aimed to identify instances of hate speech in the images. The primary focus was on identifying images that intentionally conveyed hateful sentiments. Annotators needed to distinguish between images expressing strong disagreement without resorting to offensive language and those containing genuine elements of hate speech. This differentiation aimed to guarantee accurate labeling, ensuring that images conveying genuinely hateful sentiment through the visual content, language, or a combination of both were appropriately identified.
Hate Targets: This task required annotators to identify the targets of hate in hateful images by classifying the images into one of the four classes: Undirected, Individual, Organization, and Community. Images were labeled as Undirected when they targeted abstract topics, societal themes, or ambiguous targets such as "you", and were not directed toward any specific individuals, entities, or groups. Hateful images targeting specific people including political leaders, celebrities, or activists such as "Joe Biden" and "J.K. Rowling" were annotated as Individual. Likewise, the label Community was used for instances of images targeting broader social, ethnic, or cultural groups such as "LGBT" or "trans". Lastly, images targeting corporate entities, institutions, or similar organizations such as "Chick-fil-A" and "government" were annotated as Organization.
Stance: This task involved annotating the images into either of three distinct categories: Support, Oppose, and Neutral, determined by their context within discussions on the LGBTQ+ movement. The Support label was given to images that expressed support towards the goals of the movement, agreed with efforts in fostering equal rights for LGBTQ+ individuals, and promoted awareness for the movement’s goals. The Oppose label was given to images that conveyed disagreement with the goals of the movement, denied the problems faced by individuals who identified as LGBTQ+, and dismissed the need for equal rights and acceptance. The Neutral label was given to images that were contextually relevant to the movement but did not exhibit support or opposition towards the movement.
Humor: In this annotation task, annotators were asked to identify images showcasing humor, sarcasm, or satire related to the LGBTQ+ Pride movement. Annotators were instructed to discern the presence of humor in the images regardless of whether they presented a lighthearted perspective on serious subjects. Note that annotators were asked to annotate images based on whether the creator of the image intended for it to be humorous, and not based on whether the annotator personally found it humorous. This task aimed to capture the nuanced use of text-embedded images for comedic or satirical purposes, thereby enhancing comprehension of the varied expressions within discussions on LGBTQ+.

### v2

@@ Abstract
The complexity of text-embedded images presents a formidable challenge in machine learning given the need for multimodal understanding of multiple aspects of expression conveyed by them. While previous research in multimodal analysis has primarily focused on singular aspects such as hate speech and its subclasses, this study expands this focus to encompass multiple aspects of linguistics: hate, targets of hate, stance, and humor. We introduce a novel dataset PrideMM comprising 5,063 text-embedded images associated with the LGBTQ+ Pride movement, thereby addressing a serious gap in existing resources. We conduct extensive experimentation on PrideMM by using unimodal and multimodal baseline methods to establish benchmarks for each task. Additionally, we propose a novel framework MemeCLIP for efficient downstream learning while preserving the knowledge of the pre-trained CLIP model. The results of our experiments show that MemeCLIP achieves superior performance compared to previously proposed frameworks on two real-world datasets. We further compare the performance of MemeCLIP and zero-shot GPT-4 on the hate classification task. Finally, we discuss the shortcomings of our model by qualitatively analyzing misclassified samples. Our code and dataset are publicly available at: https://github.com/SiddhantBikram/MemeCLIP.
@@
Figure 1: Samples of text-embedded images from the PrideMM dataset annotated across four aspect labels. The labels are in the form of {Hate, Target, Stance, Humor}. For Hate, {0, 1} correspond to No Hate and Hate respectively. For Target, {0, 1, 2, 3} correspond to hate targeted towards Undirected, Individual, Community, and Organization respectively. For Stance, {0, 1, 2} correspond to Neutral, Support, and Oppose respectively. For Humor, {0, 1} correspond to No Humor and Humor respectively.
@@ 1 Introduction
However, with opinions being expressed freely, hate speech becomes prevalent, often directed towards individuals, organizations, and even marginalized communities Thapa et al. (2023), targeting them with vitriol and prejudice Lingiardi et al. (2020); Imperato et al. (2023). Particularly, the LGBTQ+ movement stands as a prominent subject of online discourse, where memes serve as vehicles of both solidarity and resistance, reflecting the multifaceted dynamics of attitudes and perceptions within the community and beyond Gal et al. (2016). In this context, the distinction between humor and harm becomes blurred, as memes straddle the line between satire and offense, challenging researchers and platforms alike to navigate the complexities of online content moderation Langvardt (2017). Previous attempts that endeavored to suppress such content have resulted in the discriminative suppression of all LGBTQ+ content Griffin (2022, 2024), which can harm the awareness and acceptance of this community. Thus, understanding the nuances of hate speech, opinions, and intended humor within memes becomes paramount for fostering an inclusive digital environment and combating online discrimination.
@@
The analysis of text-embedded images is particularly challenging given the need for contextual understanding and the prevalence of ambiguity and subjectivity in them Sherratt (2022). Accordingly, the multi-aspect nature of our dataset provides a more holistic view of the diverse themes usually expressed through memes. Through PrideMM, we aim to cultivate a more profound understanding of interactions on social media through memes and facilitate the development of multimodal content moderation methods to make the internet a safer space. We implement a range of baseline and state-of-the-art hate speech detection models to establish benchmarks for each task of PrideMM. Sample images from PrideMM are illustrated in Figure 1 alongside their annotation labels.
We further propose MemeCLIP, a novel framework that leverages the knowledge of the Contrastive Language-Image Pre-Training (CLIP) model Radford et al. (2021) by using multiple lightweight modules for multimodal and multi-aspect meme classification. We employ linear layers to effectively disentangle image and text representations in CLIP’s multimodal embedding space. We utilize Feature Adapters to preserve the prior knowledge of CLIP and adapt its embedding spaces to the meme classification task while avoiding overfitting on smaller datasets. We further implement a cosine classifier alongside Semantic-Aware initialization Shi et al. (2023) to make it more robust to the class imbalances that may exist in datasets such as PrideMM and HarMeme Pramanick et al. (2021a) that are representative of real-world data distributions. Distinct from previous multimodal meme classification frameworks, MemeCLIP is trained end-to-end in a single step and does not rely on extraneous models to create augmented data. Our main contributions can be summarized as follows:
@@
  • We benchmark PrideMM by using various unimodal and multimodal methods including existing multimodal frameworks proposed for meme classification.
@@ 2.1 Multimodal Datasets
Multimodal image-text analysis has seen significant strides in recent years owing to the widespread popularity and availability of image-text pairs across social media. With the increasing need for hate speech and offensive content detection, multimodal datasets for hate speech detection have seen a particular surge. One of the first datasets in this domain was the Hateful Meme Challenge (HMC) dataset Kiela et al. (2020), containing synthetic memes designed to convey contrastive implications from the image and text modalities that target religion, race, disability, and sex. Similarly, the Harm-C Pramanick et al. (2021a) and Harm-P Pramanick et al. (2021b) datasets comprise memes related to the COVID-19 pandemic and US politics respectively that were annotated across three degrees of harmfulness and four subclasses of hate speech targets. Bhandari et al. (2023) annotated samples for hate speech detection and target classification similarly, collecting text-embedded images related to the Russia-Ukraine conflict from Twitter, Facebook (FB), and Reddit. Tangentially, Suryawanshi et al. (2020) employed extensive annotation guidelines to create the MultiOFF dataset for offensive content detection, consisting of memes collected from Reddit, Facebook, Twitter, and Instagram. In an effort to discern humor often expressed in memes, Tanaka et al. (2022) created a humor detection dataset by proposing a pipeline to extract memes devoid of interpersonal influence on the perception of humor. To identify disinformative memes, Qu et al. (2022) introduced the DisinfoMeme dataset that contains memes related to COVID-19, the Black Lives Matter (BLM) movement, and Veganism.
@@ 2.2 Multi-aspect Datasets
Online discourse on socio-political events is often imbued with a series of human emotions, leading researchers to study the numerous aspects of linguistics expressed in them. Dacon et al. (2022) used comments collected from RedditBias Barikeri et al. (2021) related to LGBTQ+ individuals and annotated each comment for the presence of Toxicity, Severe Toxicity, Obscene, Threat, Insults, and Identity Attacks. Similarly, Gautam et al. (2020) curated a dataset of tweets related to the #MeToo movement in social media by annotating the tweets across five different aspects. Taking multi-aspect datasets one step further, Ousidhoum et al. (2019) compiled an extensive multi-aspect Twitter dataset with English, French, and Arabic samples, with each annotated for different aspects including hate and offensiveness. Table 1 provides a detailed comparison of the datasets cited in this section.
Multi-aspect data helps better encompass the spectrum of human emotions that may be associated with social media interactions. Most multimodal datasets, while only focusing on a single aspect and its sub-classes, fail to encompass the complex dynamics of emotions expressed by the masses. Our work aims to address this gap by presenting a multimodal and multi-aspect dataset comprising three different aspects- hate, topical stance, and humor, and one subclass within hate: targets of hateful speech, to enable more nuanced studies of multimodal meme data through computational methods.
@@ 2.3 Multimodal Frameworks
Recent developments in large vision-language models have incited a wave of research in methods to tackle hate speech in text-embedded images. MOMENTA Pramanick et al. (2021b) was one of the first frameworks proposed to incorporate CLIP’s vision and language encoders for multimodal hate speech classification. It extracts regions of interest from image data and named entities from text data to combine them with CLIP representations by using cross-modal attention fusion. Similarly, HateCLIPper Kumar and Nandakumar (2022) was proposed to better model cross-modal interactions between CLIP representations. Textual inversion Gal et al. (2022) has been used to integrate visual cues in the text representation space in frameworks such as ISSUES Burbi et al. (2023). Recent works make use of image caption models to extract text captions from images and learn a single language processing model Cao et al. (2023b, a). However, rather than relying on augmented data from extraneous models, our proposed framework MemeCLIP leverages the knowledge learned by CLIP’s encoders during its pre-training step to process the rich multimodal information inside each image. Additionally, we use Feature Adapters alongside residual connections to prevent overfitting as annotated datasets for multimodal meme classification generally lack a high number of samples. We further utilize a cosine classifier to make MemeCLIP more robust to imbalanced data classes, which is prevalent in multi-label tasks in this domain.
@@ 3 Dataset
In this section, we describe various aspects of our dataset including data collection, annotation guidelines, and dataset statistics. Our dataset comprises 5,063 text-embedded images that encompass memes, posters, and infographics relevant to the LGBTQ+ movement. We only include images from 2020-2024 as this period saw an upsurge of social media content in this domain Oz et al. (2023). This also allows our dataset to represent contemporary social media interactions through memes. Note that by the term LGBTQ+, we refer to all gender identities and sexual orientations inclusively.
@@ 3.1 Data Collection
To maintain diversity in the dataset, we collected data from three popular social media platforms: Facebook, Twitter, and Reddit, through manual search and extraction. For Twitter, we used hashtags such as #lgbt, #pride, #trans, #transrights, #nonbinary, and #genderidentity to filter images related to LGBTQ+ discussions. For Facebook, we targeted groups that frequently discussed LGBTQ+ content. Similarly, for Reddit, we identified subreddits where discussion related to LGBTQ+ was more prominent. Further, to ensure the relevance and quality of the dataset, the data collection process was subject to filtering criteria. Detailed filtering criteria for our dataset can be found in Appendix A.2. As different annotators may encounter and collect the same image, we sequentially employed two image deduplication tools: dupeGuru [^1] https://github.com/arsenetar/dupeguru and difPy [^2] https://github.com/elisemercury/Duplicate-Image-Finder, to search for duplicates and retain the highest quality image out of each batch of duplicates. We used the OCR application provided by Google Cloud Vision API [^3] https://cloud.google.com/vision/docs to extract textual data from the images. We removed non-alphanumeric elements such as special characters, hyperlinks, symbols, and non-English characters to reduce noisy text data and ensure data quality. Note that the text may occasionally contain unintentional noisy artifacts.
@@ 3.2 Data Annotation
We engaged five experienced annotators well-versed in NLP and computational linguistics to annotate data samples for PrideMM. The annotators had a prior understanding of the LGBTQ+ movement and meme archetypes on social media. We presented them with comprehensive annotation guidelines to ensure uniform and unbiased annotations, and asked them to annotate each image separately for all four tasks. A 3-phase annotation schema was used to ensure accurate and consistent annotations. First, a dry run was conducted to evaluate the understanding of the annotation guidelines among the annotators where every annotator was given an identical batch of 50 images for annotation. Second, a revision phase was conducted where every annotator was given another identical batch of 200 images and received a revised set of instructions based on the results of the first phase. Finally, in the consolidation phase, the annotators annotated a final batch of 50 images while discussing and revising the annotation guidelines until a consensus was reached. These steps were taken to minimize misannotations and noisy labels in the PrideMM dataset. The meticulously devised annotation guidelines were followed to ensure consistency in the annotations. Each image in our dataset was independently annotated for the three aspects and one sub-class, apart from the connection between ’Hate’ and ’Hate Targets’.
@@ 3.3 Annotation Guidelines
Hate Speech.  This task aimed to identify instances of hate speech in the images. The primary focus was on identifying images that intentionally conveyed hateful sentiments. Annotators needed to distinguish between images expressing strong disagreement without resorting to offensive language and those containing genuine elements of hate speech. This differentiation aimed to guarantee accurate labeling, ensuring that images conveying genuinely hateful sentiment through visual content, language, or a combination of both were appropriately identified.
Hate Targets.  This task required annotators to identify the targets of hate in hateful images by classifying the images into one of the four classes: Undirected, Individual, Community, and Organization. Images were labeled as Undirected when they targeted abstract topics, societal themes, or ambiguous targets like ‘you’ that were not directed toward any specific individuals, entities, or groups. Hateful images targeting specific people including political leaders, celebrities, or activists like ‘Joe Biden’ and ‘J.K. Rowling’ were annotated as Individual. Likewise, the label Community was used for instances of images targeting broader social, ethnic, or cultural groups like ‘LGBT’ or ‘trans’. Lastly, images targeting corporate entities, institutions, or similar organizations like ‘Chick-fil-A’ and ‘government" were annotated as Organization.
Stance.  This task involved annotating the images into either of three distinct categories: Support, Oppose, and Neutral, determined by their stance within the context of the LGBTQ+ movement. The Support label was given to images that expressed support towards the goals of the movement, agreed with efforts in fostering equal rights for LGBTQ+ individuals, and promoted awareness for the movement’s goals. The Oppose label was given to images that conveyed disagreement with the goals of the movement, denied the problems faced by individuals who identified as LGBTQ+, and dismissed the need for equal rights and acceptance. The Neutral label was given to images that were contextually relevant to the movement but did not exhibit support or opposition towards the movement.
Humor.  In this task, annotators were asked to identify images showcasing humor, sarcasm, or satire related to the LGBTQ+ Pride movement. Annotators were instructed to discern the presence of humor in the images regardless of whether they presented a lighthearted or insensitive perspective on serious subjects. Note that annotators were asked to annotate images based on whether the creator of the image intended for it to be humorous, and not based on whether the annotator personally found it humorous. This task aimed to capture the nuanced use of text-embedded images for comedic or satirical purposes, thereby helping disentangle hate and humor in the images related to this movement.